OpenStudy (anonymous):

lim x->0^+ (sinx)/(x+sqrootx)

terenzreignz (terenzreignz):

There's always L'hôpital for your everyday indeterminate needs...

OpenStudy (anonymous):

I you don't know L'Hopital's rule yet: For \(x\) near 0, you have \(\sin x\approx x\). \[\lim_{x\to0^+}\frac{\sin x}{x+\sqrt x}=\lim_{x\to0^+}\frac{x}{x+\sqrt x}=\lim_{x\to0^+}\frac{1}{1+x^{-1/2}}\] Multiply numerator and denominator by \(\sqrt x\): \[\lim_{x\to0^+}\frac{1}{1+x^{-1/2}}\cdot\frac{\sqrt x}{\sqrt x}=\lim_{x\to0^+}\frac{\sqrt x}{\sqrt x+1}\]